# Development of a comprehensive PMT optical model for the JUNO experiment

### v1

@@ 4.4 Comparison of the average reflection and PDE
Figure 10: The average reflectance (top) and PDE without the contributions from transmitted photons (bottom) as a function of the AOI for the HPK PMT (left), NNVT normal-QE PMT (middle) and NNVT high-QE PMT (right). The blue solid lines represent the results from the non-uniform ARC and PC, while the orange dashed lines indicate the results from the uniform ARC and PC, and the green solid lines show their ratio.
@@ 5 Conclusion
A precise PMT optical model is essential for improving the accuracy of the energy response of the JUNO detector, thereby enhancing the sensitivity and precision of JUNO’s physics outputs. In this work, we developed a comprehensive PMT optical model based on previous research, aiming at extracting some basic optical parameters of PMTs and then precisely simulating the non-uniformity of PDE across the PMT surface and variations in PDE among different PMTs. Reflectance data were collected for 669 PMTs, which were utilized to constrain the thickness map of the PC and ARC for three types of 20-inch PMTs used in JUNO: HPK PMTs, NNVT normal-QE PMTs, and NNVT high-QE PMTs, along with PDE data gathered from both the scanning station system and the container systems used for PMT mass testing.
By performing a joint fit of both the reflectance and PDE data, we obtained the thickness map of the PC and ARC for each PMT. Taking into account the optical processes within the PMTs by simulation, the CE was determined as a function of zenith angle. The refractive index and extinction coefficient of the PC and ARC, as well as the escape factor, was estimated over an extended wavelength range from 300 nm to 700 nm, covering the entire range of interest for JUNO.
Using the new parameters for each PMT, we calculated the average reflectance and PDE, and compared these results with those from the simplified PMT optical model. The results show that the PDE undergoes changes at the level of a few percent, and the reflectance exhibits significant changes of more than 30%. The comprehensive model will be incorporated into the JUNO simulation program, and its effects on the detector’s energy response will also be thoroughly evaluated. Furthermore, additional refinements to the model are anticipated once real JUNO data become available for comparison.

### v2

@@ 4.4 Comparison of the average reflectance and PDE
Figure 10: The average reflectance (top) and PDE without the contributions from the transmitted photons (bottom) as a function of the AOI for the HPK PMT (left), NNVT normal-QE PMT (middle) and NNVT high-QE PMT (right). The blue solid lines represent the results from the non-uniform ARC and PC, while the orange dashed lines indicate the results from the uniform ARC and PC, and the green solid lines show their ratio.
@@ 5 Conclusion
A precise PMT optical model is essential for improving the accuracy of the energy response of the JUNO detector, thereby enhancing the sensitivity and precision of JUNO’s physics outputs. In this study, we developed a comprehensive PMT optical model based on the previous research, aiming at extracting some basic optical parameters of PMTs and then precisely simulating the non-uniformity of the PDE across the PMT surface and the variations in the PDE among different PMTs. The reflectance data collected from 669 PMTs in this work as well as the PDE data gathered from both the scanning station system and the container system were utilized to constrain the thickness map of the PC and ARC for three types of 20-inch PMTs in JUNO: the HPK PMTs, NNVT normal-QE PMTs, and NNVT high-QE PMTs.
By performing a joint fit of both the reflectance and PDE data, we constructed the thickness map of the PC and ARC for each PMT. Taking into account the optical processes within the PMTs by simulation, the CE was determined as a function of the zenith angle by decomposing the PDE data. The refractive index and extinction coefficient of the PC and ARC, as well as the escape factor, were estimated over an extended wavelength range from 300 nm to 700 nm, covering the entire range of interest for JUNO.
Utilizing the new parameters for each PMT, we calculated the average reflectance and PDE, and compared these results with those from the simplified PMT optical model. The results show that the PDE undergoes changes at the level of a few percent, and the reflectance exhibits significant changes of more than 30%. The comprehensive model will be incorporated into the JUNO simulation program, and its effects on the detector’s energy response will also be thoroughly evaluated. Furthermore, additional refinements to the model are anticipated once real JUNO data become available for comparison.